A horizontal force of 0.5N acts on a body of mass 0.2kg. there is frictional force of 0.2 opposing the first force. What acceleration will be produced?
1
Expert's answer
2020-03-31T09:26:37-0400

The net force on a body

"F_{net}=F-F_f=0.5\:\rm N-0.2\: N=0.3\: N."

The acceleration of the body

"a=\frac{F_{net}}{m}=\frac{0.3\:\rm N}{0.2\:\rm kg}=1.5\:\rm N."